OpenStudy (anonymous):

Simplify (15/ sqr20)-(sqr20 /4)

ganeshie8 (ganeshie8):

change both fractions to same denominator

ganeshie8 (ganeshie8):

\[\dfrac{15}{\sqrt{20}} - \dfrac{\sqrt{20}}{4}\] \[\dfrac{4*15}{4*\sqrt{20}} - \dfrac{\sqrt{20}*\sqrt{20}}{4*\sqrt{20}}\] \[\dfrac{60}{4*\sqrt{20}} - \dfrac{20}{4*\sqrt{20}}\] \[\dfrac{60-20}{4*\sqrt{20}} \]

ganeshie8 (ganeshie8):

you can simplify further

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got sqr 5

ganeshie8 (ganeshie8):

\(\sqrt{5}\) is \(\large \color{Red}{\checkmark}\)

OpenStudy (anonymous):

cool, thanks, I have another question I need help with, I'll post it :)

ganeshie8 (ganeshie8):

u wlc :)